Mr Kumah is three times as old as his son,20year ago,his son was quarter his fathers age.find the sum of their ages
Answer by stanbon(75887)

Mr Kumah is three times as old as his son,20year ago,his son was quarter his fathers age.find the sum of their ages
----
Equations:
K = 3s
s-20 = (1/4)(K-20)
-------
4s-80 = 3s-20
s = 60 (son's age)
K = 3s = 180 (father's age)------ impossible
----
Sum of their ages:: 180+60 = 240 ---- also impossible
------
Please check your post for accuracy.
Cheers,
Stan H.
----------